OpenStudy (anonymous):

Is it possible for this: ¶(dy/dx) = f'(x).......¶ is integrate symbol ¶ dy = ¶ f'(x) dx My friend said she copied this from teacher's note, i don understand this, is this really possible or there is sum mistake?

OpenStudy (anonymous):

\[\frac{dy}{dx} = f'(x)\]\[dy=f'(x) dx\]\[\int dy=\int f'(x) dx\]\[y=f(x)+c\]